# Semantic Integrity Constraints: Declarative Guardrails for AI-Augmented Data Processing Systems

### v2

@@ Abstract.
AI-augmented data processing systems (DPSs) integrate large language models (LLMs) into query pipelines, allowing powerful semantic operations on structured and unstructured data. However, the reliability (a.k.a. trust) of these systems is fundamentally challenged by the potential for LLMs to produce errors, limiting their adoption in critical domains. To help address this reliability bottleneck, we introduce semantic integrity constraints (SICs)—a declarative abstraction for specifying and enforcing correctness conditions over LLM outputs in semantic queries. SICs generalize traditional database integrity constraints to semantic settings, supporting common types of constraints, such as grounding, soundness, and exclusion, with both proactive and reactive enforcement strategies.
We argue that SICs provide a foundation for building reliable and auditable AI-augmented data systems. Specifically, we present a system design for integrating SICs into query planning and runtime execution and discuss its realization in AI-augmented DPSs. To guide and evaluate the vision, we outline several design goals—covering criteria around expressiveness, runtime semantics, integration, performance, and enterprise-scale applicability—and discuss how our framework addresses each, along with open research challenges.
@@ 1. Introduction
Large language models (LLMs) have transformed the field of data management in the last couple of years, extending traditional systems with semantic processing capabilities over structured and unstructured data. In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025). We broadly refer to these systems as AI-augmented data processing systems (DPSs). While AI-augmented DPSs introduce novel ways to process data, their use of LLMs lead to potential reliability issues that manifest into phenomena like hallucinations. These reliability issues hinder the adoption of AI-augmented DPSs in high-stakes sectors, such as healthcare, law, and finance. Without guardrails, it will be difficult to fully adopt these systems.
LLM constraints—also known as assertions or guardrails—have recently been actively explored in academia and industry (Beurer-Kellner et al., 2023; Singhvi et al., 2024; Rebedea et al., 2023; Shankar et al., 2024a, b; Dong et al., 2024a; Liu et al., 2024; Shankar et al., 2024c; AI, 2023a, b; Index, 2023; TruLens, 2021; Zheng et al., 2023). These constraints leverage user-defined functions (UDFs) or LLM validators to check an LLM’s output, evaluating code or natural language predicates, respectively. A related approach is constrained decoding, where constraints are applied during the LLM’s generation to ensure that the output adheres to a specified structure (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023). While these approaches have proven their utility in enhancing LLM reliability, they are imperative and fragmented, limiting their usability.
We propose semantic integrity constraints (SICs) as a declarative, unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs. Unlike traditional integrity constraints that prevent invalid modifications to the database, SICs protect against erroneous outputs from semantic operators. Similar to their traditional counterparts, SICs make defining constraints for predominant use cases easy and more advanced cases possible, aligning with recent work that argues for specifications on LLM-based system components (Stoica et al., 2024).
@@
| Performance Awareness | Enforcement strategy is selected based on cost and reliability across proactive or reactive methods (Section 4.2) |
@@
This paper presents our vision for SICs and sketches a general framework for integrating them into AI-augmented DPSs. To evaluate the feasibility and impact of this vision, we define a set of design goals and success criteria: (1) expressiveness, to capture a broad range of common constraint types; (2) executable semantics, with well-defined runtime behavior; (3) system integration, via seamless support within data flow and query processing pipelines; (4) performance-aware enforcement, balancing reliability and cost; and (5) enterprise-scale applicability, including observability, constraint reuse, conflict detection, and auditability. These criteria, outlined in Table 1, were defined through our partnership with the Rhode Island Hospital. They guide our design and are reflected in the structure of the paper. Section 3.1 introduces the SIC interface, while Section 3.2 categorizes SICs across key enforcement classes aligned with real-world LLM use cases. Section 4 explores enforcement strategies, spanning logical and physical planning, and outlines cost-aware mechanisms. Section 5 discusses observability components. Finally, Section 6 presents open challenges in large-scale constraint management.
2. System Model
@@
3. Constraint Specification
We suggest defining SICs declaratively, allowing users to focus on constraint logic while the optimizer determines the best execution method. We outline the main components of our proposed interface and detail classes of constraints relevant to semantic data processing. User surveys should be conducted to ensure usability.
3.1. Interface
To provide a unified interface for queries and SICs, users declare constraints in the query language. Specifically, we present a relational interface that extends pipe SQL syntax (Shute et al., 2024) with semantic operators (e.g., map, aggregate, filter, join, top-) and SICs. Though, SICs are compatible with any declarative AI-augmented DPS. The interface allows users to specify two types of information to the DPS for each constraint: (1) what the constraint is and (2) how the DPS should react when the constraint is violated.
Declaring Constraints. SICs specify predicates either on generated tuple attributes from semantic operators or on the semantic operators themselves. We make this distinction because while some operators (e.g., map, aggregate) generate attributes, other operators (e.g., filter, join, top-) do not; yet, constraints should still be defined in either case. The syntax for declaring a constraint on an attribute involves a SQL expression starting with ASSERT <attribute>, followed by a keyword that refers to the constraint class (Section 3.2) and any necessary parameters. A constraint on an operator is similar: ASSERT <operator>. Users can also combine constraints via logical connectives (e.g., ASSERT <pred1> AND <pred2>).
Retry Thresholds. Users can define a retry threshold for each operator. When a constraint associated with the operator (either on the operator or on an attribute generated by the operator) is violated, the DPS retries the operator until the output satisfies the constraint or the retry threshold is reached. To specify a retry threshold, users write RETRY <threshold> after the operator.
Failure Modes. Failure modes (CONTINUE, IGNORE, or ABORT) are also associated with operators and specify what the DPS should do when a failing output reaches the retry threshold. With CONTINUE, the DPS continues processing the query, allowing errors to propagate downstream. IGNORE specifies that the DPS should ignore invalid tuples in the processing (including dependent tuples, such as those that would be in the same aggregation window). Lastly, ABORT aborts the entire query. To specify a failure mode, the user writes <mode> ON FAIL after the retry threshold.
End-To-End Example. LABEL:lst:query shows an example semantic query with various SICs. The query is declared over the ehr_table (line 1), a relation of electronic health records (EHRs). Each tuple in the relation corresponds to a patient, and we assume that it contains attributes for the patient’s ehr and dob. The query first canonicalizes the dob attribute’s value into YYYY-MM-DD (line 2) because the value could have been in another format. The expression is defined using a string prefixed with the letter “p”, which declares a prompt string. Prompt strings specify natural language functions or predicates over attributes and are evaluated using LLMs. Line 4 computes the patient’s age using standard SQL functions. Lines 5-7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history. Lastly, the query filters tuples by selecting patients likely to have sepsis (lines 10-11). Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the query (line 16).
The first constraint is declared on line 3, and the remaining constraints are specified on line 12 and onward. If the summarization step for med_hist_sum (line 8) fails its constraints (lines 13-15), the DPS retries the operator once and then CONTINUEs to propagate the tuple even if it still fails the constraints (line 9). Since no retry thresholds and failure modes are defined in the rest of the query, we assume that some default setting is used. We explain each constraint in more detail in the next section.
Semantic query with pipe SQL syntax and SICs.
@@ 3.2. Constraint Classes
Domain Constraints. In many queries, users want the generated attributes to belong to a specified domain. This includes use cases such as enforcing that a value is of a certain data type or structure, from a set of values, or less than a certain length. For example, one may check whether the predicted prescription dose from a semantic map is within a range of possible values. We extend traditional domain constraints to semantic data processing. In LABEL:lst:query, lines 3 and 15 specify domain constraints, restricting the dob to the regex representation of YYYY-MM-DD and the summary’s length to less than 1000 characters, respectively. The type definitions (e.g., STRING) of generated attributes implicitly define domain constraints. This class of SICs comprises the majority of constraint use cases (Liu et al., 2024) and is enforced deterministically.
@@
Inclusion and Exclusion (IE) Constraints. Users sometimes expect that an attribute generated by an operator includes or excludes certain terms, content, topics, or abstract features such as tones or styles. For example, given a semantic map that summarizes an EHR, users may want to exclude all patient PII from the summary while including the names of attending physicians. IE constraints address this class of predicates. In the example query, line 14 defines an exclusion constraint on med_hist_sum, which excludes test results from the medical history summary. When an IE constraint is defined with a prompt string (as in the example), LLMs are used to check for a semantic match. Users can also specify IE constraints using strings, regular expressions, or even a collection of them (e.g., a set of toxic words to exclude from the output). In such cases, deterministic methods are used to evaluate matches.
@@
Grounding Constraints. Semantic operators, such as maps or aggregates, can extract data or summarize from their inputs. However, since LLMs can hallucinate, users want to ensure that the generated data is grounded in source data. Grounding constraints enable users to enforce such invariants on generated attributes. For example, line 12 ensures that the extracted phys_exam and lab_res are exactly contained in the source EHR. Line 13 asserts that the summarized med_hist_sum is factually consistent with the source EHR. In particular, the last grounding constraint implicitly defines a grounding constraint on med_hist, since med_hist_sum is derived from med_hist, which, in turn, is derived from ehr. Formally, if a grounding constraint is defined on an attribute, the DPS will recursively enforce grounding constraints on all ancestor attributes in its lineage. Thus, each constraint only needs to compare its operator’s input and output to ensure grounding in the source data. A grounding constraint is implemented in one of two ways, depending on whether the specified attribute is generated via an extractive (map) or abstractive (map or aggregate) operator. In the extractive case, the constraint is implemented deterministically and ensures that the operator’s output is precisely contained in its input. In the abstractive case, the constraint is implemented stochastically (e.g., with an LLM or a fact-checking model (Tang et al., 2024)) and ensures that the operator’s summarized output is factually consistent with the input. Users can annotate relevant operators with the EXTRACTIVE (lines 5-7) or ABSTRACTIVE (line 8) keywords to inform the DPS of their semantics; otherwise, the DPS infers these annotations by reasoning about the operator’s prompt string with an LLM.
@@
| Grounding | Value is grounded in the source values | Extracted test results or summaries are factual w.r.t. original EHRs |
@@
Soundness Constraints. To check whether the reasoning of operators themselves is logically sound, users can declare soundness constraints. Although applicable to all operators, this constraint class is particularly useful for operators that do not generate attributes, such as filter, join, or top-. To enforce soundness constraints, the DPS uses LLMs to check the operator’s reasoning process by modifying the operator to output its reasoning steps, e.g., chain-of-thought (CoT) (Wei et al., 2022). Line 16 declares a soundness constraint on the sepsis filter. Assume that the query earlier retrieved relevant sepsis-related medical information (omitted for brevity). This retrieved information is considered to describe the symptoms of sepsis and the diagnostic rationale. Under the hood, the DPS prompts the filter’s LLM to return its CoT in addition to a boolean value that represents the selection result. The CoT first lists the premises, i.e., the patient’s medical condition and the retrieved medical information relevant to sepsis. The premises are followed by the reasoning steps that lead to the selection result. Then, the input and output of the filter’s LLM are given to an LLM validator (or several of them) to ensure that the premises are grounded in the input and that the reasoning steps are valid. The DPS can also use more sophisticated CoT prompting and verification techniques, which use structured natural language formats or programming languages (Ling et al., 2023; Lyu et al., 2023; Chen et al., 2023).
Relevance Constraints. These constraints validate if an attribute’s value is relevant for the given operator’s task; i.e., they verify whether the underlying LLM followed the instructions prompted by the user. Since relevance constraints are always appropriate, the DPS can enforce them on all generated attributes by default. We assume that this is the case in the example query. For example, the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than something irrelevant, such as their social history. These constraints can be evaluated by passing the operator’s prompt, input, and output to an LLM judge.
Assertions. Like assertions in traditional integrity constraints, these constraints enable users to specify arbitrary predicates that must hold in the data flow. The DPS cannot offer many optimizations for assertions due to their black-box nature; however, such constraints give users the flexibility to express conditions that are not supported by the primitives described previously.
@@ 4. Constraint Enforcement
Constraint enforcement consists of two phases: detection and recovery. First, a constraint detects any violations in the operator’s output. Upon detection of a violation, the operator attempts to recover. This distinction aids in reasoning about optimizing constraint enforcement, as discussed in this section.

### v3

@@ Abstract.
AI-augmented data processing systems (DPSs) integrate large language models (LLMs) into query pipelines, allowing powerful semantic operations on structured and unstructured data. However, the reliability (a.k.a. trust) of these systems is fundamentally challenged by the potential for LLMs to produce errors, limiting their adoption in critical domains. To help address this reliability bottleneck, we introduce semantic integrity constraints (SICs)—a declarative abstraction for specifying and enforcing correctness conditions over LLM outputs in semantic queries. SICs generalize traditional database integrity constraints to semantic settings, supporting common types of constraints, such as grounding, soundness, and exclusion, with both reactive and proactive enforcement strategies.
We argue that SICs provide a foundation for building reliable and auditable AI-augmented data systems. Specifically, we present a system design for integrating SICs into query planning and runtime execution and discuss its realization in AI-augmented DPSs. To guide and evaluate our vision, we outline several design goals—covering criteria around expressiveness, runtime semantics, integration, performance, and enterprise-scale applicability—and discuss how our framework addresses each, along with open research challenges.
@@ 1. Introduction
Large language models (LLMs) have transformed the field of data management in the last couple of years, extending traditional systems with semantic processing capabilities over structured and unstructured data. In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025; Agarwal and Huang, 2025; Madden et al., 2024). We broadly refer to these systems as AI-augmented data processing systems (DPSs). While AI-augmented DPSs introduce novel ways to process data, their use of LLMs lead to potential reliability issues that manifest into phenomena like hallucinations. These reliability issues hinder the adoption of AI-augmented DPSs in high-stakes sectors, such as healthcare, law, and finance. Without guardrails, it will be difficult to adopt these systems fully.
LLM assertions have actively been explored for offline evaluations (TruLens, 2021; Zheng et al., 2023; Shankar et al., 2024b, c; AI, 2023b; Index, 2023) and runtime guardrails (AI, 2023a; Rebedea et al., 2023; Dong et al., 2024a; Singhvi et al., 2024; Liu et al., 2024; Loula et al., 2025). These assertions invoke custom code, ML classifiers, or LLM judges to check an LLM’s output with respect to various criteria. A related technique is constrained decoding, where constraints are applied during the LLM’s generation to ensure that the output adheres to a specified structure (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023). While these approaches have proven their utility in enhancing LLM reliability, they are imperative and fragmented, limiting their usability.
To address these limitations, we propose semantic integrity constraints (SICs) as a declarative, unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs. Unlike traditional integrity constraints that prevent invalid modifications to the database, SICs protect against erroneous outputs from semantic operators. Similar to their traditional counterparts, SICs make defining constraints for predominant use cases easy and more advanced cases possible, aligning with recent work that argues for specifications on LLM-based system components (Stoica et al., 2024).
@@
| Performance Awareness | Enforcement strategy is selected based on cost and reliability across reactive and proactive methods (Section 4.2) |
@@
This paper presents our vision for SICs and sketches a general framework for integrating them into AI-augmented DPSs. To evaluate the feasibility and impact of this vision, we define a set of design goals and success criteria: (1) expressiveness, to capture a broad range of common constraint types; (2) executable semantics, with well-defined runtime behavior; (3) system integration, via seamless support within data flow and query processing pipelines; (4) performance-aware enforcement, balancing reliability and cost; and (5) enterprise-scale applicability, including observability, constraint reuse, conflict detection, and auditability. These criteria, outlined in Table 1, were defined through our partnership with the Rhode Island Hospital. They guide our design and are reflected in the structure of the paper. Section 2 outlines the proposed system architecture. Section 3 introduces the SIC interface and categorizes SICs across key enforcement classes aligned with real-world LLM use cases. Section 4 explores enforcement strategies, spanning logical and physical planning, and outlines cost-aware mechanisms. Section 5 discusses observability components. Finally, Section 6 presents open challenges in large-scale constraint management.
2. System Architecture
@@
3. Constraint Specification
We suggest defining SICs declaratively, allowing users to focus on constraint logic while the optimizer determines the best execution method. We outline the main components of our proposed interface and detail classes of constraints relevant to semantic data processing. Additional surveys should be conducted to ensure usability.
3.1. Interface
To provide a unified interface for queries and SICs, users declare constraints in the query language. We present a relational interface that extends pipe SQL syntax (Shute et al., 2024) with semantic operators (e.g., map, aggregate, filter, join, top-) and SICs; however, SICs are compatible with any declarative AI-augmented DPS.
Declaring Constraints. SICs specify predicates either on generated tuple attributes from semantic operators or on the semantic operators themselves. We make this distinction because while some operators (e.g., map, aggregate) generate attributes, other operators (e.g., filter, join, top-) do not; yet, constraints should still be defined in either case. The syntax for declaring a constraint on an attribute involves a SQL expression starting with ASSERT, followed by a predicate involving the attribute’s name, the type of constraint (Section 3.2), and any necessary parameters. A constraint on an operator is similar. Users can also combine constraints via logical connectives (e.g., ASSERT <pred1> AND <pred2>).
Retry Thresholds. Users can define a retry threshold for each constraint. When a constraint associated with the operator (either on the operator or on an attribute generated by the operator) is violated, the DPS retries the operator until the output satisfies the constraint or the retry threshold is reached. To specify a retry threshold, users write RETRY <threshold> after the constraint.
Failure Modes. Failure modes (CONTINUE, IGNORE, or ABORT) are also associated with constraints and specify what the DPS should do when a failing output reaches the retry threshold. With CONTINUE, the DPS continues processing the query, allowing errors to propagate downstream. IGNORE specifies that the DPS should ignore invalid tuples in the processing (including dependent tuples, such as those that would be in the same aggregation window). Lastly, ABORT cancels the entire query. To specify a failure mode, the user writes <mode> ON FAIL after the retry threshold.
End-To-End Example. LABEL:lst:query shows an example semantic query with various SICs. The query is declared over the ehr_table (line 1), a relation of electronic health records (EHRs). Each tuple in the relation corresponds to a patient, and we assume that it contains attributes for the patient’s ehr and dob. The query first canonicalizes the dob attribute’s value into YYYY-MM-DD (line 2) because the value could have been in another format. The expression involves a string prefixed with the letter “p”, which defines a prompt string. Prompt strings can specify natural language functions or predicates over attributes and are evaluated using LLMs. Line 4 computes the patient’s age using standard SQL functions. Lines 5-7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history. Lastly, the query filters tuples by selecting patients likely to have sepsis (line 9). Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the query (line 14).
The first constraint is declared on line 3, and the remaining constraints are specified on line 10 and onward. If the summarization step for med_hist_sum (line 8) fails the constraint on line 12, the DPS retries the operator once and then CONTINUEs to propagate the tuple even if it still fails the constraint again (line 13). Since no retry thresholds and failure modes are defined in the rest of the query, we assume that some default setting is used. We explain each constraint in more detail in the next section.
Semantic query with pipe SQL syntax and SICs.
@@ 3.2. Constraint Classes
Domain Constraints. In many queries, users want generated attributes to belong to a specified domain. This includes use cases such as enforcing that a value conforms to a certain data type or structure, belongs to a set of values, or is less than a certain length. For example, one may check whether the predicted prescription dose from a semantic map is within a range of possible values. We extend traditional domain constraints to semantic data processing. In LABEL:lst:query, line 3 and the second conjunct on line 11 specify domain constraints, restricting the dob to the regex for YYYY-MM-DD and the summary’s length to less than 1000 characters, respectively. Type declarations (e.g., STRING) of generated attributes implicitly define domain constraints. This class of SICs comprises the majority of constraint use cases (Liu et al., 2024) and is enforced deterministically.
@@
Inclusion and Exclusion (IE) Constraints. Users sometimes expect that an attribute generated by an operator includes or excludes certain terms, content, topics, or abstract features such as tones or styles. For example, given a semantic map that summarizes an EHR, users may want to exclude all the patient’s PII from the summary while including the names of attending physicians. IE constraints address this class of predicates. Line 12 defines an exclusion constraint on med_hist_sum, which excludes test results from the medical history summary. When an IE constraint is defined with a prompt string (as in the example), LLMs are used to check for a semantic match. Users can also specify IE constraints using strings, regexes, or even a collection of them (e.g., a set of toxic words to exclude from the output). In such cases, deterministic methods are used to evaluate matches.
@@
Grounding Constraints. Semantic operators, such as maps or aggregates, can extract data or summarize from their inputs. However, since LLMs can hallucinate, users want to ensure that the generated data is grounded in source data. Grounding constraints enforce such invariants. For example, line 10 ensures that the extracted phys_exam and lab_res are exactly contained in the source EHR, while the first conjunct on line 11 asserts that the summarized med_hist_sum is factually consistent with the source EHR. In particular, the last grounding constraint implicitly defines a grounding constraint on med_hist, since med_hist_sum is derived from med_hist, which, in turn, is derived from ehr. Formally, if a grounding constraint is defined on an attribute, the DPS will recursively enforce grounding constraints on all ancestor attributes in its lineage. Thus, each constraint only needs to compare its operator’s input and output to ensure grounding in the source data. A grounding constraint is implemented in one of two ways, depending on whether the specified attribute is generated via an extractive (map) or abstractive (map or aggregate) operator. In the extractive case, the constraint is implemented deterministically and ensures that the operator’s output is precisely contained in its input. In the abstractive case, the constraint is implemented stochastically (e.g., with an LLM or a fact-checking model (Tang et al., 2024)) and ensures that the operator’s summarized output is factually consistent with the input. Users can annotate relevant operators with the EXTRACTIVE (lines 5-7) or ABSTRACTIVE (line 8) keywords to inform the DPS of their semantics; otherwise, the DPS infers these annotations by reasoning about the operator’s prompt string with an LLM.
@@
| Grounding | Value is grounded in the source values | Extracted test results or summaries are factual w.r.t. source EHRs |
@@
Soundness Constraints. To check whether the reasoning of an operator is logically sound, users can declare soundness constraints. Although applicable to all operators, this constraint class is particularly useful for operators that do not generate attributes, such as filter, join, or top-. To enforce soundness constraints, the DPS uses LLMs to check the operator’s reasoning, e.g., chain-of-thought (CoT) (Wei et al., 2022). As an example, line 14 declares a soundness constraint on the sepsis filter. Assume that the query earlier retrieved relevant sepsis-related medical information (omitted for brevity). This retrieved information is considered to describe the symptoms of sepsis and the diagnostic rationale. Under the hood, the DPS prompts the filter’s LLM to return its CoT in addition to a boolean value that represents the selection result. The CoT first lists the premises, i.e., the patient’s medical condition and the retrieved medical information relevant to sepsis. The premises are followed by the reasoning steps that lead to the selection result. Then, the input and output of the filter’s LLM are given to an LLM judge (or several of them) to ensure that the premises are grounded in the input and that the reasoning steps are valid. The DPS can also use more sophisticated CoT prompting and verification techniques, which leverage structured natural language formats (Ling et al., 2023).
Relevance Constraints. These constraints validate if an attribute’s value is relevant for the given operator’s task; i.e., they verify whether the underlying LLM followed the instructions prompted by the user. Since relevance constraints are always appropriate, the DPS can enforce them on all generated attributes by default. We assume that this is the case in the example query. For instance, the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than something irrelevant, such as their social history. These constraints can be evaluated by passing the operator’s prompt, input, and output to an LLM judge.
Assertions. Like assertions in traditional integrity constraints, this class of constraints enables users to specify arbitrary predicates that must hold in the data flow. The DPS cannot offer many optimizations for assertions due to their black-box nature; however, such constraints give users the flexibility to express conditions that are not supported by the classes described previously.
@@ 4. Constraint Enforcement
Constraint enforcement consists of two phases: detection and recovery. First, a constraint detects any violations in the operator’s output. If a violation is detected, the operator attempts to recover. This distinction helps us reason about optimizing constraint enforcement, as discussed in this section.